A Bibliography of Supercomputing ’2011

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Abstract

This bibliography records articles presented at the Supercomputing ’2011 conference.

Title word cross-reference

1, 2, 3 [SMDS11]. 2 [YIM11, YBPH11].

-D [YIM11].

1.44 [LBKF11]. 10-petaflops [YWY^{+}11]. 100G [CL11]. ’11 [ACM11].

2.0 [SAT^{+}11]. 2011 [Hua11, LCK11].

abstraction [CHE11]. accelerated [MNSM11]. acceleration [VHMP11, WQY^{+}11]. access [CSS11]. accurate [CHE11, KKO^{+}11]. Adaptive [JHN11, PLP^{+}11]. Address [PWL^{+}11]. Advanced [YWY^{+}11]. aggregated [HLD11]. AICS [YWY^{+}11]. algebra [SBD11]. algorithm [ZBS11]. algorithms [BJLR11]. allocation [HM11, PSLJ11]. Amazon [ZLZ^{+}11]. Analysis [LCK11, BHK11, CS11, CHA^{+}11]. analytics [LJA^{+}11]. application [GIM^{+}11, KKO^{+}11, MH11]. Applications [PWL^{+}11, BFR11, CH11, IHIR11, NC11, SK11, SSWC11, ZLZ^{+}11, ZDJ11]. approach [SVC^{+}11]. architecture [BFR11, KD11]. architectures [HGD11, SSWC11]. archiver [Car11]. array [BWL^{+}11]. array-based [BWL^{+}11]. atmospheric [TAT^{+}11]. Atomistic [LBKF11]. atoms [HIT^{+}11, MSZ^{+}11]. Auto [MH11, WOCS11]. Auto-scaling [MH11]. auto-tuning [WOCS11]. AutomaDeD [LGdS^{+}11]. automatic [GJ11]. Avoiding [BJGK11]. aware [CH11, HLD11, PSLJ11, SBD11, VHMP11].

balancer [SK11]. bandwidth [BCK11]. Barnes [ZBS11]. based [BWL^{+}11, DJP^{+}11, KKO^{+}11, SVC^{+}11]. batch [KPB^{+}11]. Benchmark [KGL^{+}11]. Best [MSBM11]. between [RKHK11]. beyond [BCJ^{+}11]. biofluidics [BBE^{+}11]. biomolecular [MSZ^{+}11]. BlobCR [NC11]. blood [GIM^{+}11]. Blue [CEH^{+}11, VHMP11]. Boltzmann [WOCS11]. brain [GIM^{+}11]. breadth [BM11]. breadth-first [BM11]. building [DJP^{+}11].

cache [FPRK11, SVC^{+}11]. cache-friendly [SVC^{+}11]. caching [FPRK11]. calculations [HIT^{+}11, WWJ^{+}11]. capability [Har11]. car [JHN11]. catalog [WMLB11]. centers [BC11]. Challenges [AFL^{+}11, BC11]. checkpoint [NC11]. checkpoint-restart [NC11]. Checkpointing [BCR^{+}11]. class [WR11]. Cloud [ZLZ^{+}11, MH11, PSLJ11]. clouds [BRE11, LBZ^{+}11, NC11]. cluster [HM11, KKO^{+}11, ZLZ^{+}11]. clusters [MSBM11, PG11, RKHK11, WWJ^{+}11]. co [KDS^{+}11, TPO^{+}11]. co-design [KDS^{+}11]. co-designing [TPO^{+}11]. code [IJ11, MMK^{+}11]. codebase [SSWC11]. collectives [SBD11]. Communication [PWL^{+}11, SBD11]. Community [WMC^{+}11]. complex [LPAZ11]. compositing [MKPH11]. compressed [LJA^{+}11]. Computational [YWY^{+}11, GIM^{+}11, KPB^{+}11, PJR^{+}11]. computations [MNSM11]. compute [ZLZ^{+}11]. computer [HIT^{+}11]. Computing [LCK11, YWY^{+}11, Ben11, BC11, GIC^{+}11, HM11, LBZ^{+}11, LCH^{+}11]. concurrent [FPRK11]. condensed [HLD11]. Conference [LCK11]. configuration [BRE11]. consumption [GBL^{+}11]. ’cool [SK11]. cooling [CEM11]. coordination [SYS^{+}11]. Copernicus [PLP^{+}11]. copy [IHIR11]. core [BBE^{+}11, CHE11, IJ11, SVC^{+}11, YIM11]. correlated [YIM11]. cost [LBZ^{+}11, MH11]. coupled [KKO^{+}11]. CPU [MMK^{+}11]. CudaDMA [BCK11].

D [YIM11, YBPH11]. data

[CH11, CHA^{+}11, KPB^{+}11, LJA^{+}11, PSHI11, TSP^{+}11, VHMP11]. data-intensive [CH11, KPB^{+}11, TSP^{+}11]. database [WMLB11]. datacenters [GBL^{+}11]. deadlines [MH11]. debugging [LGdS^{+}11]. Deep [Har11]. Defense [Ben11]. dendritic [SAT^{+}11]. dense [NTDD11, SBD11]. density [WWJ^{+}11, YIM11]. Department [Ben11]. dependencies [LFG^{+}11]. deployment [MSBM11]. design [FHL^{+}11, KDS^{+}11, TSP^{+}11, YIM11]. designing [TPO^{+}11]. detection [CL11, PSHI11]. device [LBKF11]. DGEMM [TLT^{+}11]. direct [BJGK11]. disk [NC11]. display [WUE^{+}11]. distance [RKHK11]. distributed [BFR11, BM11, HGD11, PSHI11, WMLB11, WOCS11]. domain [DJP^{+}11, KWA^{+}11]. driven [LJA^{+}11, MSZ^{+}11]. duality [TSP^{+}11]. Dymaxion [CSS11]. dynamic [FPRK11]. dynamics [PLP^{+}11].

early [BHK11]. Earth [WMC^{+}11, TAT^{+}11]. easy [SMDS11]. effective [DEG^{+}11]. efficiency [FHL^{+}11]. Efficient [PSHI11, KDS^{+}11, NC11]. eigensolver [YBPH11]. eigenvalue [HLD11]. electricity [LBZ^{+}11]. electron [HIT^{+}11]. Enabling [MSZ^{+}11]. End [SKY11, ZDJ11]. End-to-end [SKY11]. energy [GBL^{+}11, KDS^{+}11, LPAZ11]. energy-efficient [KDS^{+}11]. engineering [LBKF11]. environments [CH11, WUE^{+}11]. Eulerian [IJ11]. Evaluating [FSL^{+}11, ZLZ^{+}11]. evaluation [KPB^{+}11]. evaluations [IJ11]. event [CS11]. evolution [NG11]. exascale [FSL^{+}11, LCH^{+}11]. excited [KKO^{+}11]. excited-state [KKO^{+}11]. exploring [CHE11]. Extracting [WOCS11]. extreme [LJA^{+}11]. extreme-scale [LJA^{+}11].

facilities [TPO^{+}11]. factorization [BJLR11]. failures [HKG^{+}11]. Fast [TLT^{+}11, HGD11, NG11]. fault [BGTK^{+}11]. Fermi [TLT^{+}11]. field [SAT^{+}11]. file [Car11, SYS^{+}11, TSP^{+}11, WR11]. fine [HLD11]. fine-grained [HLD11]. First [HIT^{+}11, BM11]. First-principles [HIT^{+}11]. Flexible [HM11, SKY11]. floating [DEG^{+}11]. floating-point [DEG^{+}11]. flow [GIM^{+}11, JHN11]. forms [HLD11]. framework [GJ11]. free [YBPH11]. frequency [LFG^{+}11]. friendly [SVC^{+}11]. FTI [BGTK^{+}11]. full [JHN11]. functional [WWJ^{+}11]. Fusion [PWL^{+}11].

GAMER [SSWC11]. Gene [CEH^{+}11, VHMP11]. Gene/P [VHMP11]. Gene/Q [CEH^{+}11]. generation [GJ11]. genomic [MBK11]. Global [PWL^{+}11, TAT^{+}11]. GPU [BCK11, MNSM11, MMK^{+}11, TLT^{+}11, WWJ^{+}11]. GPU-accelerated [MNSM11]. GPUs [BCJ^{+}11, NTDD11]. grained [HLD11]. graph [YBPH11]. graphs [YBPH11]. green [GBL^{+}11]. GreenSlot [GBL^{+}11]. grids [BRE11]. GROPHECY [MMK^{+}11]. group [YIM11]. GT5D [IJ11]. Gyrokinetic [MIW^{+}11, PWL^{+}11, IJ11].

Hadoop [BWL^{+}11, WQY^{+}11]. Hardware [KDS^{+}11]. Hardware/software [KDS^{+}11]. hashing [GLJ11]. HDFS [TSP^{+}11]. heterogeneous [CSS11, HKG^{+}11, HGD11]. hierarchical [WOCS11]. High [FHL^{+}11, LCK11, SVC^{+}11, YWY^{+}11, BGTK^{+}11, Ben11, BC11, GIC^{+}11, KKO^{+}11, LBZ^{+}11, WUE^{+}11]. High-efficiency [FHL^{+}11]. high-level [KKO^{+}11]. High-Performance [YWY^{+}11, SVC^{+}11, BC11]. high-resolution [WUE^{+}11]. highest [TAT^{+}11]. Highly [MBK11]. Hot [CEM11, BJGK11]. hot-spots [BJGK11]. house [ZLZ^{+}11]. HPC [AFL^{+}11, BHK11, Har11, MIW^{+}11, MSBM11, NC11, PG11, TPO^{+}11]. Hut [ZBS11]. hybrid [BGTK^{+}11, SVC^{+}11].

I/O [FPRK11, KPB^{+}11, SYS^{+}11, VHMP11, ZDJ11]. I/O-intensive [ZDJ11]. IaaS [NC11]. IBM [CEH^{+}11]. identification [MBK11]. IH [BHK11]. image [MKPH11, NC11]. implementation [TLT^{+}11]. implementations [KKO^{+}11]. implications [LCH^{+}11]. implicitly [MNSM11]. improve [DEG^{+}11]. Improving [SBD11]. in-house [ZLZ^{+}11]. index [CHA^{+}11]. InfiniBand [RKHK11]. initio [MBK11]. instances [ZLZ^{+}11]. Institute [YWY^{+}11, YWY^{+}11]. Integrating [WUE^{+}11]. intensive [CH11, IHIR11, KPB^{+}11, TSP^{+}11, ZDJ11]. interconnection [CEH^{+}11, RKHK11]. interface [BGTK^{+}11]. interference [CH11]. interference-aware [CH11]. International [LCK11]. Intrusion [CL11]. ISABELA [LJA^{+}11]. ISABELA-compressed [LJA^{+}11]. ISABELA-QA [LJA^{+}11].

Janus [TPO^{+}11]. Japanese [YWY^{+}11]. jobs [BCR^{+}11].

kernels [HLD11]. Keynote [Hua11].

language [DJP^{+}11]. LANL [PG11]. Large [LGdS^{+}11, WWJ^{+}11, CHA^{+}11, HKG^{+}11, LFG^{+}11, MNSM11, YBPH11]. large-scale [LFG^{+}11, MNSM11]. lattice [BCJ^{+}11, SVC^{+}11, VSZ^{+}11, WOCS11]. leading [MIW^{+}11]. level [BJGK11, CHE11, DEG^{+}11, KKO^{+}11]. levitated [WQY^{+}11]. linear [SBD11]. Liszt [DJP^{+}11]. live [IHIR11]. load [SK11]. locality [PSLJ11]. locality-aware [PSLJ11]. log [Car11]. logging [CS11]. LOGJAM [Car11]. long [RKHK11]. long-distance [RKHK11]. look [BFR11]. look-up [BFR11]. LSST [WMLB11].

machine [LBZ^{+}11]. machines [MSZ^{+}11]. MAMPO [GJ11]. management [BC11, FPRK11, MSBM11]. manycore [MIW^{+}11]. MapReduce [BFR11, PSLJ11]. massively [IJ11, SSWC11]. matrix [NTDD11, YIM11]. matrix-vector [NTDD11]. MAximum [GJ11]. measure [KGL^{+}11, Kra11, LFG^{+}11]. meet [MH11]. memory [BCK11, BM11, CSS11, GLJ11, HLD11, IHIR11, LCH^{+}11, PSHI11, WR11]. memory-aware [HLD11]. merge [WQY^{+}11]. mesh [DJP^{+}11]. mesh-based [DJP^{+}11]. message [CEH^{+}11, MSZ^{+}11]. message-driven [MSZ^{+}11]. methods [HGD11, KKO^{+}11]. metrics [Har11]. migration [IHIR11]. million [BBE^{+}11, MSZ^{+}11]. million-core [BBE^{+}11]. minimize [MH11]. mining [BFR11]. model [JHN11, MNSM11, TAT^{+}11, WMC^{+}11]. Modeling [HKG^{+}11, HGKS11, KDS^{+}11, NG11]. modernization [Ben11]. molecular [PLP^{+}11]. monitoring [AFL^{+}11, Ben11, DEG^{+}11]. motif [MBK11]. movement [VHMP11]. MPI [SVC^{+}11, ZLZ^{+}11]. Multi [SSWC11, CHE11, IJ11, MIW^{+}11, SVC^{+}11, WUE^{+}11, YIM11]. multi- [MIW^{+}11]. multi-core [CHE11, IJ11, SVC^{+}11, YIM11]. Multi-science [SSWC11]. multi-touch [WUE^{+}11]. Multicore [GJ11, MSZ^{+}11]. multicore-optimized [MSZ^{+}11]. multiplication [NTDD11]. multipole [HGD11]. multiscale [GIM^{+}11, VSZ^{+}11].

Multithreaded [PWL^{+}11, GJ11].

nanoelectronic [LBKF11]. nanowire [HIT^{+}11]. National [YWY^{+}11]. network [CEH^{+}11, SKY11, WQY^{+}11]. Networking [LCK11]. networks [BJGK11]. nothing [WMLB11]. November [LCK11]. numbers [SMDS11]. NWSC [KGL^{+}11].

O [FPRK11, KPB^{+}11, SYS^{+}11, VHMP11]. O-intensive [ZDJ11]. One [GIC^{+}11]. operational [PG11]. optimization [LPAZ11]. Optimized [IHIR11, MSZ^{+}11]. Optimizing [NTDD11, ZBS11, BCK11, CSS11]. organizational [BC11].

P [VHMP11]. paradigm [GIM^{+}11, PLP^{+}11]. Parallel [BM11, CHA^{+}11, HLD11, SMDS11, BCR^{+}11, CHE11, HKG^{+}11, IJ11, KWA^{+}11, LGdS^{+}11, MNSM11, PSHI11, PLP^{+}11, SK11, SSWC11, SVC^{+}11, SYS^{+}11]. Parallelization [YIM11]. partitioning [YBPH11]. past [JHN11]. patterns [CSS11]. PDE [DJP^{+}11]. software [KDS^{+}11]. Performance [HGKS11, IJ11, LCK11, WMC^{+}11, YWY^{+}11, BHK11, BGTK^{+}11, Ben11, BC11, DEG^{+}11, GIC^{+}11, KGL^{+}11, Kra11, LBZ^{+}11, MMK^{+}11, PG11, SVC^{+}11, SBD11, TAT^{+}11, WOCS11]. performances [LBKF11]. perspective [BC11]. Peta [SAT^{+}11]. Peta-scale [SAT^{+}11]. Petaflop [BBE^{+}11]. petaflops [YWY^{+}11]. petascale [MSZ^{+}11]. PFlop [LBKF11]. PFlop/s [LBKF11]. phase [SAT^{+}11]. phase-field [SAT^{+}11]. phenomena [VSZ^{+}11]. Physis [MNSM11]. placement [LBZ^{+}11]. plane [WWJ^{+}11]. Platforms [PWL^{+}11, IJ11, YIM11]. point [DEG^{+}11]. populations [NG11]. porphyrin [KKO^{+}11]. porphyrin-based [KKO^{+}11]. portable [DJP^{+}11]. POwer [GJ11]. POWER7 [BHK11]. POWER7-IH [BHK11]. Practice [ACM11, PJR^{+}11]. practices [MSBM11]. pre [IHIR11]. pre-copy [IHIR11]. principles [HIT^{+}11]. problems [HLD11]. Proceedings [LCK11]. process [FSL^{+}11]. processing [BWL^{+}11]. production [MSBM11, RKHK11]. program [Ben11]. programming [MNSM11]. programs [PSHI11]. project [Har11, KD11]. projection [MMK^{+}11]. pseudopotential [WWJ^{+}11]. Purlieus [PSLJ11]. PVFS [TSP^{+}11].

Q [CEH^{+}11]. QA [LJA^{+}11]. QCD [BCJ^{+}11, SVC^{+}11]. QoS [SKY11, ZDJ11]. QR [BJLR11]. Qserv [WMLB11]. quantum [VSZ^{+}11, YIM11]. qubit [VSZ^{+}11]. queries [KPB^{+}11]. query [BWL^{+}11, CHA^{+}11, LJA^{+}11]. query-driven [LJA^{+}11].

race [PSHI11]. random [SMDS11]. reconciling [TSP^{+}11]. Reducing [LBZ^{+}11]. reduction [HLD11]. reliability [FSL^{+}11, LCH^{+}11]. reliable [HM11]. renormalization [YIM11]. replication [FSL^{+}11]. Reports [ACM11]. requests [SKY11]. Research [YWY^{+}11]. reservation [SKY11]. resolution [TAT^{+}11, WUE^{+}11]. resource [HM11, Har11, PSLJ11, SKY11]. restart [NC11]. running [ZLZ^{+}11]. runtime [MSZ^{+}11].

s [LBKF11]. SC’11 [LCK11]. Scalable [GLJ11, HGD11, KKO^{+}11, LPAZ11, BRE11, Car11, CHE11, MBK11, YBPH11]. Scale [PWL^{+}11, CHA^{+}11, LGdS^{+}11, LJA^{+}11, LFG^{+}11, MNSM11, MKPH11, SAT^{+}11, WWJ^{+}11, WOCS11, YBPH11]. scale-free [YBPH11]. Scaling [BCJ^{+}11, MH11, MSZ^{+}11]. scheduling [CH11, GBL^{+}11, SKY11]. Science [YWY^{+}11, PJR^{+}11, SSWC11]. scientific [KGL^{+}11, LJA^{+}11]. SciHadoop [BWL^{+}11].

SCMFS [WR11]. search [BM11]. Seattle [LCK11]. seismic [KDS^{+}11]. semantic [BRE11]. Server [SYS^{+}11, FHL^{+}11]. Server-side [SYS^{+}11]. service [BRE11]. shared [GLJ11, WMLB11, ZDJ11]. shared-nothing [WMLB11]. side [SYS^{+}11]. silicon [HIT^{+}11]. similarity [LFG^{+}11]. Simplified [KWA^{+}11]. simulation [CHE11, JHN11, SAT^{+}11]. simulations [BBE^{+}11, GIM^{+}11, MIW^{+}11, MSZ^{+}11, VSZ^{+}11]. Simulator [TAT^{+}11]. single [SSWC11]. skeletons [MMK^{+}11]. snapshots [NC11]. Sniper [CHE11]. SNL [GIC^{+}11]. solidification [SAT^{+}11]. solution [MKPH11]. solver [NG11]. solvers [DJP^{+}11]. Space [PWL^{+}11]. specialization [BCK11]. specific [DJP^{+}11]. SPOTlight [PG11]. spots [BJGK11]. stability [PG11]. staging [VHMP11]. State [ACM11, KKO^{+}11]. states [HIT^{+}11]. stencil [MNSM11]. stochastic [LPAZ11]. stop [GIC^{+}11]. Storage [LCK11, FPRK11, WR11, ZDJ11]. strategies [BCR^{+}11]. streaming [KPB^{+}11]. strongly [YIM11]. strongly-correlated [YIM11]. Suite [KGL^{+}11]. supercomputer [KGL^{+}11, SAT^{+}11, YWY^{+}11]. supercomputers [GLJ11, MNSM11]. supercomputing [VHMP11]. support [GIC^{+}11, ZDJ11]. survey [PJR^{+}11]. Sustained [Ben11, Kra11, LBKF11]. symmetric [HLD11, NTDD11]. synthetic [GJ11]. System [DEG^{+}11, LCH^{+}11, WMC^{+}11, BBE^{+}11, TSP^{+}11, WR11]. System-level [DEG^{+}11]. systematic [HGKS11]. systems [BHK11, BGTK^{+}11, Ben11, BM11, CSS11, FSL^{+}11, GJ11, HM11, HKG^{+}11, KKO^{+}11, LPAZ11, MIW^{+}11, SVC^{+}11, SYS^{+}11, TPO^{+}11, VHMP11, YIM11, ZDJ11].

tasks [LGdS^{+}11]. Techniques [PWL^{+}11]. technology [KD11]. testing [PG11]. text [BFR11]. theories [KKO^{+}11]. theory [WWJ^{+}11]. threaded [SVC^{+}11]. threaded-MPI [SVC^{+}11]. throughput [KGL^{+}11]. Tiled [BJLR11]. time [LFG^{+}11]. tolerance [BGTK^{+}11]. tolerating [HKG^{+}11]. toolkit [CS11]. TOP500 [KD11]. Topology [VHMP11, SBD11]. Topology-aware [VHMP11]. toroidal [MIW^{+}11]. touch [WUE^{+}11]. trace [KD11]. TRACON [CH11]. traversal [KWA^{+}11]. trends [KD11]. TSUBAME [SAT^{+}11]. tuning [HGKS11, WOCS11]. turbulence [KPB^{+}11, VSZ^{+}11]. turbulent [JHN11]. two [BBE^{+}11, BJGK11, RKHK11]. two-level [BJGK11].

U.S. [Ben11]. Ultra [PWL^{+}11, WOCS11]. Ultra-Scale [PWL^{+}11, WOCS11]. unified [Car11]. unit [CEH^{+}11]. Unitary [VSZ^{+}11]. UPC [ZBS11]. usage [Har11]. use [RKHK11]. useful [Kra11]. user [GIC^{+}11]. users [ZDJ11]. Using [KD11, BFR11, HLD11, KGL^{+}11, NC11, SVC^{+}11, YBPH11, ZDJ11]. utilization [DEG^{+}11].

vector [NTDD11]. versus [ZLZ^{+}11]. via [BCK11, SKY11, SBD11, WOCS11]. viability [FSL^{+}11]. Virtual [FPRK11, HM11, LBZ^{+}11, NC11]. virtualized [CH11]. virus [GJ11, NG11].

WA [LCK11]. Wallaby [BRE11]. warm [CEM11]. warp [BCK11]. water [CEM11]. wave [WWJ^{+}11]. wide [Har11]. workflows [MH11]. workloads [FPRK11, LFG^{+}11]. World [TAT^{+}11]. World-highest [TAT^{+}11].

References

ACM:2011:SSP

[ACM11] ACM, editor. SC ’11 State of the Practice Reports. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-1139-3. LCCN ????

Allcock:2011:CHM

[AFL^{+}11] William (Bill) Allcock, Evan Felix, Mike Lowe, Randal Rheinheimer, and Joshi Fullop. Challenges of HPC monitoring. In ACM [ACM11], pages 22:1–22:6. ISBN 1-4503-1139-3. LCCN ????

Bernaschi:2011:PBS

[BBE^{+}11] Massimo Bernaschi, Mauro Bisson, Toshio Endo, Satoshi Matsuoka, Massimiliano Fatica, and Simone Melchionna. Petaflop biofluidics simulations on a two million-core system. In Lathrop et al. [LCK11], pages 4:1–4:12. ISBN 1-4503-0771-X. LCCN ????

Berente:2011:CMH

[BC11] Nicholas Berente and Jennifer Claggett. Challenges in the management of high-performance computing centers: an organizational perspective. In ACM [ACM11], pages 18:1–18:8. ISBN 1-4503-1139-3. LCCN ????

Babich:2011:SLQ

[BCJ^{+}11] R. Babich, M. A. Clark, B. Joó, G. Shi, R. C. Brower, and S. Gottlieb. Scaling lattice QCD beyond 100 GPUs. In Lathrop et al. [LCK11], pages 70:1–70:11. ISBN 1-4503-0771-X. LCCN ????

Bauer:2011:COG

[BCK11] Michael Bauer, Henry Cook, and Brucek Khailany. CudaDMA: optimizing GPU memory bandwidth via warp specialization. In Lathrop et al. [LCK11], pages 12:1–12:11. ISBN 1-4503-0771-X. LCCN ????

Bougeret:2011:CSP

[BCR^{+}11] Marin Bougeret, Henri Casanova, Mikael Rabie, Yves Robert, and Frédéric Vivien. Checkpointing strategies for parallel jobs. In Lathrop et al. [LCK11], pages 33:1–33:11. ISBN 1-4503-0771-X. LCCN ????

Bennett:2011:SSP

[Ben11] Paul M. Bennett. Sustained systems performance monitoring at the U.S. Department of Defense high performance computing modernization program. In ACM [ACM11], pages 3:1–3:11. ISBN 1-4503-1139-3. LCCN ????

Balkir:2011:DLA

[BFR11] Atilla Soner Balkir, Ian Foster, and Andrey Rzhetsky. A distributed look-up architecture for text mining applications using MapReduce. In Lathrop et al. [LCK11], pages 59:1–59:11. ISBN 1-4503-0771-X. LCCN ????

Bautista-Gomez:2011:FHP

[BGTK^{+}11] Leonardo Bautista-Gomez, Seiji Tsuboi, Dimitri Komatitsch, Franck Cappello, Naoya Maruyama, and Satoshi Matsuoka. FTI: high performance fault tolerance interface for hybrid systems. In Lathrop et al. [LCK11], pages 32:1–32:12. ISBN 1-4503-0771-X. LCCN ????

Barker:2011:EPA

[BHK11] Kevin J. Barker, Adolfy Hoisie, and Darren J. Kerbyson. An early performance analysis of POWER7-IH HPC systems. In Lathrop et al. [LCK11], pages 42:1–42:11. ISBN 1-4503-0771-X. LCCN ????

Bhatele:2011:AHS

[BJGK11] Abhinav Bhatele, Nikhil Jain, William D. Gropp, and Laxmikant V. Kale. Avoiding hot-spots on two-level direct networks. In Lathrop et al. [LCK11], pages 76:1–76:11. ISBN 1-4503-0771-X. LCCN ????

Bouwmeester:2011:TQF

[BJLR11] Henricus Bouwmeester, Mathias Jacquelin, Julien Langou, and Yves Robert. Tiled QR factorization algorithms. In Lathrop et al. [LCK11], pages 7:1–7:11. ISBN 1-4503-0771-X. LCCN ????

Buluc:2011:PBF

[BM11] Aydin Buluç and Kamesh Madduri. Parallel breadth-first search on distributed memory systems. In Lathrop et al. [LCK11], pages 65:1–65:12. ISBN 1-4503-0771-X. LCCN ????

Benton:2011:WSS

[BRE11] William C. Benton, Robert H. Rati, and Erik J. Erlandson. Wallaby: a scalable semantic configuration service for grids and clouds. In ACM [ACM11], pages 10:1–10:10. ISBN 1-4503-1139-3. LCCN ????

Buck:2011:SAB

[BWL^{+}11] Joe B. Buck, Noah Watkins, Jeff LeFevre, Kleoni Ioannidou, Carlos Maltzahn, Neoklis Polyzotis, and Scott Brandt. Sci-Hadoop: array-based query processing in Hadoop. In Lathrop et al. [LCK11], pages 66:1–66:11. ISBN 1-4503-0771-X. LCCN ????

Cardo:2011:LSU

[Car11] Nicholas P. Cardo. LOGJAM: a scalable unified log file archiver. In ACM [ACM11], pages 23:1– 23:9. ISBN 1-4503-1139-3. LCCN ????

Chen:2011:IBG

[CEH^{+}11] Dong Chen, Noel A. Eisley, Philip Heidelberger, Robert M. Senger, Yutaka Sugawara, Sameer Kumar, Valentina Salapura, David L. Satterfield, Burkhard Steinmacher-Burow, and Jeffrey J. Parker. The IBM Blue Gene/Q interconnection network and message unit. In Lathrop et al. [LCK11], pages 26:1–26:10. ISBN 1-4503-0771-X. LCCN ????

Coles:2011:HWW

[CEM11] Henry Coles, Michael Ellsworth, and David J. Martinez. “hot” for warm water cooling. In ACM [ACM11], pages 17:1– 17:10. ISBN 1-4503-1139-3. LCCN ????

Chiang:2011:TIA

[CH11] Ron C. Chiang and H. Howie Huang. TRACON: interferenceaware scheduling for dataintensive applications in virtualized environments. In Lathrop et al. [LCK11], pages 47:1–47:12. ISBN 1-4503-0771-X. LCCN ????

Chou:2011:PIQ

[CHA^{+}11] Jerry Chou, Mark Howison, Brian Austin, Kesheng Wu, Ji Qiang, E. Wes Bethel, Arie Shoshani, Oliver Rübel, Prabhat, and Rob D. Ryne. Parallel index and query for large scale data analysis. In Lathrop et al. [LCK11], pages 30:1–30:11. ISBN 1-4503-0771-X. LCCN ????

Carlson:2011:SEL

[CHE11] Trevor E. Carlson, Wim Heirman, and Lieven Eeckhout. Sniper: exploring the level of abstraction for scalable and accurate parallel multi-core simulation. In Lathrop et al. [LCK11], pages 52:1–52:12. ISBN 1-4503-0771-X. LCCN ????

Campbell:2011:ID

[CL11] Scott Campbell and Jason Lee. Intrusion detection at 100G. In ACM [ACM11], pages 14:1–14:9. ISBN 1-4503-1139-3. LCCN ????

Carey:2011:TEA

[CS11] James Carey and Philip Sanders. A toolkit for event analysis and logging. In ACM [ACM11], pages 24:1–24:7. ISBN 1-4503-1139-3. LCCN ????

Che:2011:DOM

[CSS11] Shuai Che, Jeremy W. Sheaffer, and Kevin Skadron. Dymaxion: optimizing memory access patterns for heterogeneous systems. In Lathrop et al. [LCK11], pages 13:1–13:11. ISBN 1-4503-0771-X. LCCN ????

DelVento:2011:SLM

[DEG^{+}11] Davide Del Vento, Thomas Engel, Siddhartha S. Ghosh, David L. Hart, Rory Kelly, Si Liu, and Richard Valent. System-level monitoring of floating-point performance to improve effective system utilization. In ACM [ACM11], pages 5:1–5:6. ISBN 1-4503-1139-3. LCCN ????

DeVito:2011:LDS

[DJP^{+}11] Zachary DeVito, Niels Joubert, Francisco Palacios, Stephen Oakley, Montserrat Medina, Mike Barrientos, Erich Elsen, Frank Ham, Alex Aiken, Karthik Duraisamy, Eric Darve, Juan Alonso, and Pat Hanrahan. Liszt: a domain specific language for building portable mesh-based PDE solvers. In Lathrop et al. [LCK11], pages 9:1–9:12. ISBN 1-4503-0771-X. LCCN ????

Frachtenberg:2011:HES

[FHL^{+}11] Eitan Frachtenberg, Ali Heydari, Harry Li, Amir Michael, Jacob Na, Avery Nisbet, and Pierluigi Sarti. High-efficiency server design. In Lathrop et al. [LCK11], pages 27:1–27:11. ISBN 1-4503-0771-X. LCCN ????

Frasca:2011:VCD

[FPRK11] Michael Frasca, Ramya Prabhakar, Padma Raghavan, and Mahmut Kandemir. Virtual I/O caching: dynamic storage cache management for concurrent workloads. In Lathrop et al. [LCK11], pages 38:1–38:11. ISBN 1-4503-0771-X. LCCN ????

Ferreira:2011:EVP

[FSL^{+}11] Kurt Ferreira, Jon Stearley, James H. Laros III, Ron Oldfield, Kevin Pedretti, Ron Brightwell, Rolf Riesen, Patrick G. Bridges, and Dorian Arnold. Evaluating the viability of process replication reliability for exascale systems. In Lathrop et al. [LCK11], pages 44:1–44:12. ISBN 1-4503-0771-X. LCCN ????

Goiri:2011:GSE

[GBL^{+}11] Íñigo Goiri, Ryan Beauchea, Kien Le, Thu D. Nguyen, Md. E. Haque, Jordi Guitart, Jordi Torres, and Ricardo Bianchini. GreenSlot: scheduling energy consumption in green datacenters. In Lathrop et al. [LCK11], pages 20:1–20:11. ISBN 1-4503-0771-X. LCCN ????

Greenfield:2011:OSH

[GIC^{+}11] John A. Greenfield, Lisa G. Ice, Sophia E. Corwell, Karen Haskell, Constantine Pavlakos, and John P. Noe. One stop high performance computing user support at SNL. In ACM [ACM11], pages 26:1–26:6. ISBN 1-4503-1139-3. LCCN ????

Grinberg:2011:NCP

[GIM^{+}11] Leopold Grinberg, Joseph A. Insley, Vitali Morozov, Michael E. Papka, George Em Karniadakis, Dmitry Fedosov, and Kalyan Kumaran. A new computational paradigm in multiscale simulations: application to brain blood flow. In Lathrop et al. [LCK11], pages 5:1–5:12. ISBN 1-4503-0771-X. LCCN ????

Ganesan:2011:MMP

[GJ11] Karthik Ganesan and Lizy K.

John. MAximum Multicore POwer (MAMPO): an automatic multithreaded synthetic power virus generation framework for multicore systems. In Lathrop et al. [LCK11], pages 53:1–53:12. ISBN 1-4503-0771-X. LCCN ????

Goodman:2011:SHS

[GLJ11] Eric Goodman, M. Nicole Lemaster, and Edward Jimenez. Scalable hashing for shared memory supercomputers. In Lathrop et al. [LCK11], pages 41:1–41:11. ISBN 1-4503-0771-X. LCCN ????

Hart:2011:DWM

[Har11] David Hart. Deep and wide metrics for HPC resource capability and project usage. In ACM [ACM11], pages 1:1–1:7. ISBN 1-4503-1139-3. LCCN ????

Hu:2011:SFM

[HGD11] Qi Hu, Nail A. Gumerov, and Ramani Duraiswami. Scalable fast multipole methods on distributed heterogeneous architectures. In Lathrop et al. [LCK11], pages 36:1–36:12. ISBN 1-4503-0771-X. LCCN ????

Hoefler:2011:PMS

[HGKS11] Torsten Hoefler, William Gropp, William Kramer, and Marc Snir. Performance modeling for systematic performance tuning. In ACM [ACM11], pages 6:1–6:12. ISBN 1-4503-1139-3. LCCN ????

Hasegawa:2011:FPC

[HIT^{+}11] Yukihiro Hasegawa, Jun-Ichi Iwata, Miwako Tsuji, Daisuke Takahashi, Atsushi Oshiyama, Kazuo Minami, Taisuke Boku, Fumiyoshi Shoji, Atsuya Uno, Motoyoshi Kurokawa, Hikaru Inoue, Ikuo Miyoshi, and Mitsuo Yokokawa. First-principles calculations of electron states of a silicon nanowire with 100,000 atoms on the K computer. In Lathrop et al. [LCK11], page 1:1. ISBN 1-4503-0771-X. LCCN ????

Heien:2011:MTH

[HKG^{+}11] Eric Heien, Derrick Kondo, Ana Gainaru, Dan LaPine, Bill Kramer, and Franck Cappello. Modeling and tolerating heterogeneous failures in large parallel systems. In Lathrop et al. [LCK11], pages 45:1–45:11. ISBN 1-4503-0771-X. LCCN ????

Haidar:2011:PRC

[HLD11] Azzam Haidar, Hatem Ltaief, and Jack Dongarra. Parallel reduction to condensed forms for symmetric eigenvalue problems using aggregated fine-grained and memory-aware kernels. In Lathrop et al. [LCK11], pages 8:1–8:11. ISBN 1-4503-0771-X. LCCN ????

Hacker:2011:FRA

[HM11] Thomas J. Hacker and Kanak Mahadik. Flexible resource allocation for reliable virtual cluster computing systems. In Lathrop et al. [LCK11], pages 48:1–48:12. ISBN 1-4503-0771-X. LCCN ????

Huang:2011:SK

[Hua11]  Jen-Hsun Huang.  SC 2011

keynote. In Lathrop et al. [LCK11], page ?? ISBN 1-4503-0771-X. LCCN ????

Ibrahim:2011:OPC

[IHIR11] Khaled Z. Ibrahim, Steven Hofmeyr, Costin Iancu, and Eric Roman. Optimized pre-copy live migration for memory intensive applications. In Lathrop et al. [LCK11], pages 40:1–40:11. ISBN 1-4503-0771-X. LCCN ????

Idomura:2011:PEG

[IJ11] Yasuhiro Idomura and Sébastien Jolliet. Performance evaluations of gyrokinetic Eulerian code GT5D on massively parallel multi-core platforms. In ACM [ACM11], pages 4:1–4:9. ISBN 1-4503-1139-3. LCCN ????

Jansson:2011:AST

[JHN11] Niclas Jansson, Johan Hoffman, and Murtazo Nazarov. Adaptive simulation of turbulent flow past a full car model. In ACM [ACM11], pages 20:1–20:8. ISBN 1-4503-1139-3. LCCN ????

Kogge:2011:UTT

[KD11] Peter M. Kogge and Timothy J. Dysart. Using the TOP500 to trace and project technology and architecture trends. In Lathrop et al. [LCK11], pages 28:1–28:11. ISBN 1-4503-0771-X. LCCN ????

Krueger:2011:HSC

[KDS^{+}11] Jens Krueger, David Donofrio, John Shalf, Marghoob Mohiyuddin, Samuel Williams, Leonid Oliker, and Franz-Josef Pfreund. Hardware/software co-design for energy-efficient seismic modeling. In Lathrop et al. [LCK11], pages 73:1–73:12. ISBN 1-4503-0771-X. LCCN ????

Kelly:2011:NBS

[KGL^{+}11] Rory C. Kelly, Siddartha S. Ghosh, Si Liu, Davide Del Vento, and Richard A. Valent. The NWSC Benchmark Suite using scientific throughput to measure supercomputer performance. In ACM [ACM11], pages 7:1–7:5. ISBN 1-4503-1139-3. LCCN ????

Kowalski:2011:SIA

[KKO^{+}11] Karol Kowalski, Sriram Krishnamoorthy, Ryan M. Olson, Vinod Tipparaju, and E. Aprà. Scalable implementations of accurate excited-state coupled cluster theories: application of highlevel methods to porphyrinbased systems. In Lathrop et al. [LCK11], pages 72:1–72:10.

ISBN 1-4503-0771-X. LCCN ????

Kanov:2011:SEB

[KPB^{+}11] Kalin Kanov, Eric Perlman, Randal Burns, Yanif Ahmad, and Alexander Szalay. I/O streaming evaluation of batch queries for data-intensive computational turbulence. In Lathrop et al. [LCK11], pages 29:1–29:10. ISBN 1-4503-0771-X. LCCN ????

Kramer:2011:HMU

[Kra11] William Kramer. How to measure useful, sustained performance. In ACM [ACM11], pages 2:1–2:17. ISBN 1-4503-1139-3. LCCN ????

Kendall:2011:SPD

[KWA^{+}11] Wesley Kendall, Jingyuan Wang, Melissa Allen, Tom Peterka, Jian Huang, and David Erickson. Simplified parallel domain traversal. In Lathrop et al. [LCK11], pages 10:1–10:11. ISBN 1-4503-0771-X. LCCN ????

Luisier:2011:AND

[LBKF11] Mathieu Luisier, Timothy B.

Boykin, Gerhard Klimeck, and Wolfgang Fichtner. Atomistic nanoelectronic device engineering with sustained performances up to 1.44 PFlop/s. In Lathrop et al. [LCK11], pages 2:1–2:11. ISBN 1-4503-0771-X. LCCN ????

Le:2011:REC

[LBZ^{+}11] Kien Le, Ricardo Bianchini, Jingru Zhang, Yogesh Jaluria, Jiandong Meng, and Thu D. Nguyen. Reducing electricity cost through virtual machine placement in high performance computing clouds. In Lathrop et al. [LCK11], pages 22:1–22:12. ISBN 1-4503-0771-X. LCCN ????

Li:2011:SIM

[LCH^{+}11] Sheng Li, Ke Chen, Ming-Yu Hsieh, Naveen Muralimanohar, Chad D. Kersey, Jay B. Brockman, Arun F. Rodrigues, and Norman P. Jouppi. System implications of memory reliability in exascale computing. In Lathrop et al. [LCK11], pages 46:1–46:12. ISBN 1-4503-0771-X. LCCN ????

Lathrop:2011:SPI

[LCK11] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Lassnig:2011:SMT

[LFG^{+}11] Mario Lassnig, Thomas Fahringer, Vincent Garonne, Angelos Molfetas, and Martin Barisits. A similarity measure for time, frequency, and dependencies in large-scale workloads. In Lathrop et al. [LCK11], pages 43:1–43:11. ISBN 1-4503-0771-X. LCCN ????

Laguna:2011:LSD

[LGdS^{+}11] Ignacio Laguna, Todd Gamblin, Bronis R. de Supinski, Saurabh Bagchi, Greg Bronevetsky, Dong H. Anh, Martin Schulz, and Barry Rountree. Large scale debugging of parallel tasks with AutomaDeD. In Lathrop et al. [LCK11], pages 50:1–50:10. ISBN 1-4503-0771-X. LCCN ????

Lakshminarasimhan:2011:IQQ

[LJA^{+}11] Sriram Lakshminarasimhan, John Jenkins, Isha Arkatkar, Zhenhuan Gong, Hemanth Kolla, Seung-Hoe Ku, Stephane Ethier, Jackie Chen, C. S. Chang, Scott Klasky, Robert Latham, Robert Ross, and Nagiza F. Samatova. ISABELA-QA: querydriven analytics with ISABELAcompressed extreme-scale scientific data. In Lathrop et al. [LCK11], pages 31:1–31:11. ISBN 1-4503-0771-X. LCCN ????

Lubin:2011:SSO

[LPAZ11] Miles Lubin, Cosmin G. Petra, Mihai Anitescu, and Victor Zavala. Scalable stochastic optimization of complex energy systems. In Lathrop et al. [LCK11], pages 64:1–64:10. ISBN 1-4503-0771-X. LCCN ????

Marchand:2011:HSI

[MBK11] Benoît Marchand, Vladimir B. Bajic, and Dinesh K. Kaushik. Highly scalable ab initio genomic motif identification. In Lathrop et al. [LCK11], pages 56:1–56:10. ISBN 1-4503-0771-X. LCCN ????

Mao:2011:ASM

[MH11] Ming Mao and Marty Humphrey. Auto-scaling to minimize cost and meet application deadlines in cloud workflows. In Lathrop et al. [LCK11], pages 49:1–49:12. ISBN 1-4503-0771-X. LCCN ????

Madduri:2011:GTS

[MIW^{+}11] Kamesh Madduri, Khaled Z. Ibrahim, Samuel Williams, Eun-Jin Im, Stephane Ethier, John Shalf, and Leonid Oliker. Gyrokinetic toroidal simulations on leading multi- and manycore HPC systems. In Lathrop et al. [LCK11], pages 23:1–23:12. ISBN 1-4503-0771-X. LCCN ????

Moreland:2011:ICS

[MKPH11] Kenneth Moreland, Wesley Kendall, Tom Peterka, and Jian Huang. An image compositing solution at scale. In Lathrop et al. [LCK11], pages 25:1–25:10. ISBN 1-4503-0771-X. LCCN ????

Meng:2011:GGP

[MMK^{+}11] Jiayuan Meng, Vitali A. Morozov, Kalyan Kumaran, Venkatram Vishwanath, and Thomas D.

Uram. GROPHECY: GPU performance projection from CPU code skeletons. In Lathrop et al. [LCK11], pages 14:1–14:11. ISBN 1-4503-0771-X. LCCN ????

Maruyama:2011:PIP

[MNSM11] Naoya Maruyama, Tatsuo Nomura, Kento Sato, and Satoshi Matsuoka. Physis: an implicitly parallel programming model for stencil computations on large-scale GPU-accelerated supercomputers. In Lathrop et al. [LCK11], pages 11:1–11:12. ISBN 1-4503-0771-X. LCCN ????

McLay:2011:BPD

[MSBM11] Robert McLay, Karl W. Schulz, William L. Barth, and Tommy Minyard. Best practices for the deployment and management of production HPC clusters. In ACM [ACM11], pages 9:1–9:11. ISBN 1-4503-1139-3. LCCN ????

Mei:2011:ESB

[MSZ^{+}11] Chao Mei, Yanhua Sun, Gengbin Zheng, Eric J. Bohm, Laxmikant V. Kale, James C. Phillips, and Chris Harrison. Enabling and scaling biomolecular simulations of 100 million atoms on petascale machines with a multicore-optimized messagedriven runtime. In Lathrop et al. [LCK11], pages 61:1–61:11. ISBN 1-4503-0771-X. LCCN ????

Nicolae:2011:BEC

[NC11] Bogdan Nicolae and Franck Cappello. BlobCR: efficient checkpoint-restart for HPC applications on IaaS clouds using virtual disk image snapshots. In Lathrop et al. [LCK11], pages 34:1–34:12. ISBN 1-4503-0771-X. LCCN ????

Niederbrucker:2011:FSM

[NG11] Gerhard Niederbrucker and Wilfried N. Gansterer. A fast solver for modeling the evolution of virus populations. In Lathrop et al. [LCK11], pages 74:1–74:11. ISBN 1-4503-0771-X. LCCN ????

Nath:2011:OSD

[NTDD11] Rajib Nath, Stanimire Tomov, Tingxing “Tim” Dong, and Jack Dongarra. Optimizing symmetric dense matrix-vector multiplication on GPUs. In Lathrop et al. [LCK11], pages 6:1–6:10. ISBN 1-4503-0771-X. LCCN ????

Pedicini:2011:STS

[PG11] Georgia Pedicini and Jennifer Green. SPOTlight on testing: stability, performance and operational testing of LANL HPC clusters. In ACM [ACM11], pages 25:1–25:8. ISBN 1-4503-1139-3. LCCN ????

Prabhu:2011:SPC

[PJR^{+}11] Prakash Prabhu, Thomas B. Jablin, Arun Raman, Yun Zhang, Jialu Huang, Hanjun Kim, Nick P. Johnson, Feng Liu, Soumyadeep Ghosh, Stephen Beard, Taewook Oh, Matthew Zoufaly, David Walker, and David I. August. A survey of the practice of computational science. In ACM [ACM11], pages 19:1–19:12. ISBN 1-4503-1139-3. LCCN ????

Pronk:2011:CNP

[PLP^{+}11] Sander Pronk, Per Larsson, Iman Pouya, Gregory R. Bowman, Imran S. Haque, Kyle Beauchamp, Berk Hess, Vijay S. Pande, Peter M. Kasson, and Erik Lindahl. Copernicus: a new paradigm for parallel adaptive molecular dynamics. In Lathrop et al. [LCK11], pages 60:1–60:10. ISBN 1-4503-0771-X. LCCN ????

Park:2011:EDR

[PSHI11] Chang-Seo Park, Koushik Sen, Paul Hargrove, and Costin Iancu. Efficient data race detection for distributed memory parallel programs. In Lathrop et al. [LCK11], pages 51:1–51:12. ISBN 1-4503-0771-X. LCCN ????

Palanisamy:2011:PLA

[PSLJ11] Balaji Palanisamy, Aameek Singh, Ling Liu, and Bhushan Jain. Purlieus: locality-aware resource allocation for MapReduce in a cloud. In Lathrop et al. [LCK11], pages 58:1–58:11. ISBN 1-4503-0771-X. LCCN ????

Preissl:2011:MGA

[PWL^{+}11] Robert Preissl, Nathan Wichmann, Bill Long, John Shalf, Stephane Ethier, and Alice Koniges. Multithreaded global address space communication techniques for gyrokinetic fusion applications on ultra-scale platforms. In Lathrop et al. [LCK11], pages 12:1–12:11. ISBN 1-4503-0771-X. LCCN ????

Richling:2011:LDI

[RKHK11] Sabine Richling, Heinz Kredel, Steffen Hau, and Hans-Günther Kruse. A long-distance Infini-Band interconnection between two clusters in production use. In ACM [ACM11], pages 15:1–15:8. ISBN 1-4503-1139-3. LCCN ????

Shimokawabe:2011:PSP

[SAT^{+}11] Takashi Shimokawabe, Takayuki Aoki, Tomohiro Takaki, Toshio Endo, Akinori Yamanaka, Naoya Maruyama, Akira Nukada, and Satoshi Matsuoka. Peta-scale phase-field simulation for dendritic solidification on the TSUB-AME 2.0 supercomputer. In Lathrop et al. [LCK11], pages 3:1–3:11. ISBN 1-4503-0771-X. LCCN ????

Solomonik:2011:ICP

[SBD11] Edgar Solomonik, Abhinav Bhatele, and James Demmel. Improving communication performance in dense linear algebra via topology aware collectives. In Lathrop et al. [LCK11], pages 77:1–77:11. ISBN 1-4503-0771-X. LCCN ????

Sarood:2011:CLB

[SK11] Osman Sarood and Laxmikant V.

Kale. A ’cool’ load balancer for parallel applications. In Lathrop et al. [LCK11], pages 21:1–21:11. ISBN 1-4503-0771-X. LCCN ????

Sharma:2011:EEN

[SKY11] Sushant Sharma, Dimitrios Katramatos, and Dantong Yu. Endto-end network QoS via scheduling of flexible resource reservation requests. In Lathrop et al. [LCK11], pages 68:1–68:10. ISBN 1-4503-0771-X. LCCN ????

Salmon:2011:PRN

[SMDS11] John K. Salmon, Mark A.

Moraes, Ron O. Dror, and David E. Shaw. Parallel random numbers: as easy as 1, 2, 3. In Lathrop et al. [LCK11], pages 16:1–16:12. ISBN 1-4503-0771-X. LCCN ????

Shukla:2011:MSA

[SSWC11] Hemant Shukla, Hsi-Yu Schive, Tak-Pong Woo, and Tzihong Chiueh. Multi-science applications with single codebase — GAMER — for massively parallel architectures. In Lathrop et al. [LCK11], pages 37:1–37:11. ISBN 1-4503-0771-X. LCCN ????

Smelyanskiy:2011:HPL

[SVC^{+}11] Mikhail Smelyanskiy, Karthikeyan Vaidyanathan, Jee Choi, Bálint Joó, Jatin Chhugani, Michael A. Clark, and Pradeep Dubey.

High-performance lattice QCD for multi-core based parallel systems using a cache-friendly hybrid threaded-MPI approach. In Lathrop et al. [LCK11], pages 69:1–69:11. ISBN 1-4503-0771-X. LCCN ????

Song:2011:SSC

[SYS^{+}11] Huaiming Song, Yanlong Yin, Xian-He Sun, Rajeev Thakur, and Samuel Lang. Serverside I/O coordination for parallel file systems. In Lathrop et al. [LCK11], pages 17:1–17:11. ISBN 1-4503-0771-X. LCCN ????

Takahashi:2011:WHR

[TAT^{+}11] Keiko Takahashi, Akira Azami, Yuki Tochihara, Yoshiyuki Kubo, Ken’ichi Itakura, Koji Goto, Kenryo Kataumi, Hiroshi Takahara, Yoko Isobe, Satoru Okura, Hiromitsu Fuchigami, Jun ichi Yamamoto, Toshifumi Takei, Yoshinori Tsuda, and Kunihiko Watanabe. World-highest resolution global atmospheric model and its performance on the Earth Simulator. In ACM [ACM11], pages 21:1–21:12. ISBN 1-4503-1139-3. LCCN ????

Tan:2011:FID

[TLT^{+}11] Guangming Tan, Linchuan Li, Sean Triechle, Everett Phillips, Yungang Bao, and Ninghui Sun. Fast implementation of DGEMM on Fermi GPU. In Lathrop et al. [LCK11], pages 35:1–35:11. ISBN 1-4503-0771-X. LCCN ????

Tufo:2011:JCD

[TPO^{+}11] Henry M. Tufo, Michael K. Patterson, Michael Oberg, Matthew Woitaszek, Guy Cobb, Robert Strong, and Jim Gutowski. Janus: co-designing HPC systems and facilities. In ACM [ACM11], pages 16:1–16:9. ISBN 1-4503-1139-3. LCCN ????

Tantisiriroj:2011:DDI

[TSP^{+}11] Wittawat Tantisiriroj, Seung Woo Son, Swapnil Patil, Samuel J. Lang, Garth Gibson, and Robert B. Ross. On the duality of data-intensive file system design: reconciling HDFS and PVFS. In Lathrop et al. [LCK11], pages 67:1–67:12. ISBN 1-4503-0771-X. LCCN ????

Vishwanath:2011:TAD

[VHMP11] Venkatram Vishwanath, Mark Hereld, Vitali Morozov, and Michael E. Papka. Topologyaware data movement and staging for I/O acceleration on Blue Gene/P supercomputing systems. In Lathrop et al. [LCK11], pages 19:1–19:11. ISBN 1-4503-0771-X. LCCN ????

Vahala:2011:UQL

[VSZ^{+}11] George Vahala, Min Soe, Bo Zhang, Jeffrey Yepez, Linda Vahala, Jonathan Carter, and Sean Ziegeler. Unitary qubit lattice simulations of multiscale phenomena in quantum turbulence. In Lathrop et al. [LCK11], pages 24:1–24:11. ISBN 1-4503-0771-X. LCCN ????

Worley:2011:PCE

[WMC^{+}11] Patrick H. Worley, Arthur A. Mirin, Anthony P. Craig, Mark A. Taylor, John M. Dennis, and Mariana Vertenstein. Performance of the Community Earth System Model. In Lathrop et al. [LCK11], pages 54:1–54:11. ISBN 1-4503-0771-X. LCCN ????

Wang:2011:QDS

[WMLB11] Daniel L. Wang, Serge M. Monkewitz, Kian-Tat Lim, and Jacek Becla. Qserv: a distributed shared-nothing database for the LSST catalog. In ACM [ACM11], pages 12:1–12:11. ISBN 1-4503-1139-3. LCCN ????

Williams:2011:EUS

[WOCS11] Samuel Williams, Leonid Oliker, Jonathan Carter, and John Shalf. Extracting ultra-scale Lattice Boltzmann performance via hierarchical and distributed auto-tuning. In Lathrop et al. [LCK11], pages 55:1–55:12. ISBN 1-4503-0771-X. LCCN ????

Wang:2011:HAT

[WQY^{+}11] Yandong Wang, Xinyu Que, Weikuan Yu, Dror Goldenberg, and Dhiraj Sehgal. Hadoop acceleration through network levitated merge. In Lathrop et al. [LCK11], pages 57:1–57:10. ISBN 1-4503-0771-X. LCCN ????

Wu:2011:SFS

[WR11] Xiaojian Wu and A. L. Narasimha Reddy. SCMFS: a file system for storage class memory. In Lathrop et al. [LCK11], pages 39:1–39:11. ISBN 1-4503-0771-X. LCCN ????

Westing:2011:IMT

[WUE^{+}11] Brandt Westing, Benjamin Urick, Maria Esteva, Freddy Rojas, and Weijia Xu. Integrating multi-touch in high-resolution display environments. In ACM [ACM11], pages 8:1–8:9. ISBN 1-4503-1139-3. LCCN ????

Wang:2011:LSP

[WWJ^{+}11] Long Wang, Yue Wu, Weile Jia, Weiguo Gao, Xuebin Chi, and Lin-Wang Wang. Large scale plane wave pseudopotential density functional theory calculations on GPU clusters. In Lathrop et al. [LCK11], pages 71:1–71:10. ISBN 1-4503-0771-X. LCCN ????

Yoo:2011:SEL

[YBPH11] Andy Yoo, Allison H. Baker, Roger Pearce, and Van Emden Henson. A scalable eigensolver for large scale-free graphs using 2D graph partitioning. In Lathrop et al. [LCK11], pages 63:1–63:11. ISBN 1-4503-0771-X. LCCN ????

Yamada:2011:PDM

[YIM11] Susumu Yamada, Toshiyuki Imamura, and Masahiko Machida. Parallelization design on multicore platforms in density matrix renormalization group toward 2-D quantum stronglycorrelated systems. In Lathrop et al. [LCK11], pages 62:1–62:10.

ISBN 1-4503-0771-X. LCCN ????

Yonezawa:2011:AIC

[YWY^{+}11] Akinori Yonezawa, Tadashi Watanabe, Mitsuo Yokokawa, Mitsuhisa Sato, and Kimihiko Hirao. Advanced Institute for Computational Science (AICS): Japanese National High-Performance Computing Research Institute and its 10petaflops supercomputer “K”. In ACM [ACM11], pages 13:1–13:8. ISBN 1-4503-1139-3. LCCN ????

Zhang:2011:OBH

[ZBS11] Junchao Zhang, Babak Behzad, and Marc Snir. Optimizing the Barnes–Hut algorithm in UPC. In Lathrop et al. [LCK11], pages 75:1–75:11. ISBN 1-4503-0771-X. LCCN ????

Zhang:2011:QSE

[ZDJ11] Xuechen Zhang, Kei Davis, and Song Jiang. QoS support for end users of I/O-intensive applications using shared storage systems. In Lathrop et al. [LCK11], pages 18:1–18:12. ISBN 1-4503-0771-X. LCCN ????

Zhai:2011:CVH

[ZLZ^{+}11] Yan Zhai, Mingliang Liu, Jidong Zhai, Xiaosong Ma, and Wenguang Chen. Cloud versus in-house cluster: evaluating Amazon cluster compute instances for running MPI applications. In ACM [ACM11], pages 11:1–11:10. ISBN 1-4503-1139-3. LCCN ????